# Qualitative Failures of Image Generation Models and Their Application in Detecting Deepfakes

### v1

@@ 4.2 Some Challenging Objects
Generative models face particular challenges when it comes to generating images of objects such as clocks, Lego houses, chessboards, carpets, circuit boards, basketballs, glasses of water, dice, diagrams and tables, keyboards, and computer screens. One of the reasons for this is that these types of images contain many repeated patterns, which can be difficult for the model to accurately capture. Several examples of failed attempts to generate these objects can be seen in Figures 31 and 32. This list of challenging objects can be used to assess and compare the performance of different image generation models.
@@ 5 Conclusion and Future Work
This paper lists several qualitative indicators for identifying deepfakes. These indicators not only enable us to address the issue of fake images but also underscore the differences between generated and real-world content [7]. Furthermore, they serve as a checklist for evaluating image generation models.
It should be noted that as algorithms improve, some of these clues may become obsolete over time. However, this does not mean that these models will not make any of these mistakes in generating images. It may be necessary to use a combination of these indicators to identify deepfakes, as there is no one-size-fits-all solution.
Image generation models are becoming increasingly widespread and accessible. However, in the wrong hands, these algorithms can be used to create propaganda and other forms of fake media. In a world rife with fake news [17], we have learned not to believe everything we read. Now, we must also exercise caution when it comes to visual media. The blurring of lines between reality and fiction could transform our cultural landscape from one primarily based on truth to one characterized by artificiality and deception. As we have demonstrated with the set of cues presented here, it is possible to identify fake images. In fact, we were able to use some of these indicators to detect fake faces with high accuracy in the quiz available on whichfaceisreal.com.
While visual inspection can be a valuable tool for identifying deepfakes, it may not be sufficient to detect all types of generated images. Therefore, incorporating other methods such as machine learning algorithms or forensic analysis can offer a more thorough approach to deepfake detection. Additionally, staying up-to-date of the latest techniques and technologies in this field is crucial, as it is continually advancing.
In this study, we focused on still images. However, for videos, additional indicators beyond those outlined here, such as motion and optical flow, as well as the synchronization of lip, face, and head movements over time, can also be significant factors [3]. Educating individuals on the cues outlined in this paper may aid in combatting deepfake proliferation. It would be worthwhile to investigate whether individuals can be effectively trained to become experts in this area.

### v5

@@ 4.2 Some Challenging Objects
Generative models face particular challenges when it comes to generating images of objects such as clocks, Lego houses, chessboards, carpets, circuit boards, basketballs, glasses of water, dice, diagrams and tables, keyboards, and computer screens. One of the reasons for this is that these types of images contain many repeated patterns, which can be difficult for the model to accurately capture. Several examples of failed attempts to generate these objects can be seen in Figures 33 and 34. This list of challenging objects can be used to assess and compare the performance of different image generation models.
@@ 4.3 Memorization and Copyright
4.4 Failure modes from other studies
Certain image generation techniques may incorporate failure models to provide readers with a more comprehensive understanding of their models’ limitations. For instance, the creators of the Parti image generator [37] [^5] https://parti.research.google/ have presented some examples of such failure cases, which are illustrated in Figure 37. These failure cases can be categorized into the errors discussed earlier. It is recommended that researchers in this field consider including a discussion of their models’ failure models as a best practice.
Figure 37: Sample failure of the Parti image generation model. Please refer to here to see high resolution images.
@@ 5 Conclusion and Future Work
This paper lists several qualitative indicators for identifying generated images and deepfakes. These indicators not only enable us to address the issue of fake images but also underscore the differences between generated and real-world content [7]. Furthermore, they serve as a checklist for evaluating image generation models.
It should be noted that as algorithms improve, some of these clues may become obsolete over time. However, this does not mean that these models will not make any of these mistakes in generating images. It may be necessary to use a combination of these indicators to identify generated images, as there is no one-size-fits-all solution.
Image generation models are becoming increasingly widespread and accessible. However, in the wrong hands, these algorithms can be used to create propaganda and other forms of fake media. In a world rife with fake news [20], we have learned not to believe everything we read. Now, we must also exercise caution when it comes to visual media. The blurring of lines between reality and fiction could transform our cultural landscape from one primarily based on truth to one characterized by artificiality and deception. As we have demonstrated with the set of cues presented here, it is possible to identify fake images. In fact, in an informal investigation, we were able to use some of these indicators to detect fake faces with high accuracy in the quiz available on whichfaceisreal.com. Subsequent research can assess the extent to which these cues contribute to the detection of generated images and deepfakes by conducting behavioral experiments involving human participants.
Although visual inspection can be useful in identifying generated images, it may not be comprehensive enough to detect all types of generated content. Thus, integrating alternative approaches such as machine learning algorithms or forensic analysis can provide a more comprehensive strategy. Moreover, it is vital to stay informed about the latest advancements and techniques in this field, as it is continuously evolving.
Be aware that certain instances exist where authentic images may appear deceptive. Therefore, caution should be exercised when employing these cues to discern the authenticity of an image. Please see [8].
In this study, we focused on still images. However, for videos, additional indicators beyond those outlined here, such as motion and optical flow, as well as the synchronization of lip, face, and head movements over time, can also be significant factors [3]. One can undertake comparable initiatives to investigate indicators for identifying counterfeit audio. Educating individuals on the cues outlined in this paper may aid in combating deepfake proliferation. It would be worthwhile to investigate whether individuals can be effectively trained to become experts in this area.